OpenStudy (goformit100):

A train 125 m long passes a man, running at 5 km/hr in the same direction in which the train is going, in 10 seconds. The speed of the train is: A. 45 km/hr B. 50 km/hr C. 54 km/hr D. 55 km/hr

OpenStudy (goformit100):

@lilminiwin

OpenStudy (anonymous):

d =rt again my friend

OpenStudy (anonymous):

10s =1/360 h In 10s the man will cover 5/360 km and the train is 0.125 km long. So the total distance covered by the train in 10s is 0.125+5/360 and its speed is v=s/t v=0.125*360 + 5 =50 km/h